OpenStudy (anonymous):

how to factor out e^x from Wronskian W(e^2x,e^x)?

OpenStudy (anonymous):

e^(2x)= e^(x+x)=e^(x)*e^(x) so w=e^(x)*(e^(x),1)

OpenStudy (anonymous):

| e^(2x) e^(x) | | 2e^(2x) e^(x) | = e^(3x) - 3^(3x) = -e^(3x) e^(x) | e^(x) 1 | = e^(x) (0) - e^(x) = -e^(2x) | e^(x) 0 |